# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 4.1.2 Neural Information Squeezer (NIS)
This architectural design with invertible neural network can not only reduce the complexity of computation, but can also separate the input information from the micro-states into noise and EI for prediction clearly. The authors proved a mathematical theorem stating that, as the neural network is converged by training, the mutual information on the adjacent macro states approaches to the mutual information of the two adjacent-step micro states in the data. This means, all the information dropped by the encoder is almost irrelevant with the prediction. All the useful information is squeezed into the information bottleneck of the dynamics learner at the macro-level. And another theorem has proved that the degree of information squeezing for micro-states is increased as we reduce the dimension of the macro states, and thus more and more useless information in the original micro-state data is thrown out by the narrower and narrower information bottleneck, the macro-state dynamics learner.
To solve the problem of causal emergence identification, we first optimize the neural networks, and the dimension averaged EI, $\mathcal{J}_{q}$ can be obtained for different $q$. When $q=p$, the derived $\mathcal{J}_{p}$ is the dimension averaged EI for the micro-dynamics. Finally, the calculation of
@@
is the measure of dimension averaged causal emergence (see Equation 25. If $\Delta\mathcal{J}>0$, we can judge that causal emergence occurs.
Numeric examples are given to show the effectiveness of the whole framework. Interestingly, the neural information squeezer can work on the time series data with discrete micro-state although the whole framework is built based on continuous dynamical systems. Particularly, the model can work out the same coarse-graining strategy and macro-level dynamic as in the example of boolean network with four nodes in Hoel’s original paper, although any information about the node grouping and state mapping is not given to the framework.
It is also interesting to compare NIS with the mentioned frameworks or models in previous sections. Compare to the theory of computational mechanics, NIS can be treated as a kind of $\epsilon$-machine because any encoded macro-states can be understood as the states, and when the whole framework is well-trained such that it can predict the future micro-states precisely, the encoded macro-states converge to effective states which can be treated as causal states in computational mechanics. However, the objective that maximize the effective information (EI) in NIS has no correspondence.
NIS also has similarity with G-emergence because it also adopts the ideas of Granger causality: the effective macro-state is optimized by predicting the micro-state in the next time step. However, there are several obvious differences between the two frameworks. In theory of G-emergence, the macro-state must be selected by hand at first, however, it is optimized automatically. Second, NIS uses neural networks to predict the future states, while, G-emergence uses auto-regressive techniques to fit the data. Due to the universal approximate theorem(see  Hornik (1991); Teshima et al. (2020)), neural networks are more advantage.
NIS provides a solution for causal emergence identification from data, and it is the first time to solve the problem of the automatic construction of coarse-graining strategy and the macro-dynamics from data for causal emergence identification.
Another important benefit of NIS which is to be noticed is that, what NIS has solved actually is a generalized causal emergence identification problem but not limited on a justification. Because after training, NIS can not only identify if causal emergence occurs, but also can obtain a neural network for coarse-graining the data and a neural network to simulate macro-dynamics. This may provide us richer information than just a justification.
@@ 4.1.3 Neural Information Squeezer Plus (NIS+)
One of the biggest problem of NIS is that it does not directly solve the EI maximization problem by optimizing the neural networks (encoder, decoder, and dynamics learner). Instead, it searchs for the optimal dimension of the macro-state space $q$ such that EI can be maximized. Therefore, for given $q$, the neural networks are optimal for reconstructing the micro-dynamics (minimize Equation 43) but not for EI. Second, NIS can not support complex computations such as the coarse-graining functions containing multiple steps of information conversion and discarding.
To tackle these problems, NIS+ is proposed in ref Yang et al. (2023). Mathematically, the problem of EI maximization can be converted into a machine learning problem according to the definition of EI and the variational inequality and probability reweighting technique. After this conversion, the minimization problem can be solved by training three neural networks $\psi_{\omega},f_{\theta},g_{\theta^{\prime}}$ with parameters $\omega,\theta$ and $\theta^{\prime}$, respectively. Formally, the problem of minimization of loss function problem without constraints can be written:
@@
where $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t}))$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t+1}))$ are the macro-states. $\lambda$ is a Lagrangian multiplier which will be taken as a hyper-parameter in experiments. $w(\boldsymbol{x}_{t})$ is the inverse probability weights which is defined as:
@@
where $\tilde{p}$ is the new distribution of macro-states $\boldsymbol{y}_{t}$ after intervention for $do(\boldsymbol{y}_{t}\sim U_{q})$, and $p$ is the natural distribution of the data. In practice, $p(\boldsymbol{y}_{t})$ is estimated through kernel density estimation (KDE) Rosenblatt (1956). The approximated distribution, $\tilde{p}(\boldsymbol{y}_{t})$, is assumed to be a uniform distribution, which is characterized by a constant value. Consequently, the weight $w$ is computed as the ratio of these two distributions. Theorems are proved to guarantee the correctness of this conversion. The overall framework of NIS+ can be depicted in Figure 15.
Figure 15: The workflow and architecture of our proposed framework, Neural Information Squeezer Plus (NIS+), for discovering causal emergence within data. (a) Various forms of input data. (b) The framework of our proposed model, NIS+, which incorporates our previous model, NIS. The boxes represent functions or neural networks, and the arrow pointing to a cross represents the operation of information discarding. $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t+1}$ represent the observational data of micro-states, while $\hat{\boldsymbol{x}}_{t+1}$ represents the predicted micro-state. $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})$ represent the macro-states obtained by encoding the micro-states using the encoder. $\hat{\boldsymbol{y}}_{t}=\phi(\hat{\boldsymbol{x}}_{t})$ and $\hat{\boldsymbol{y}}_{t+1}=\phi(\hat{\boldsymbol{x}}_{t+1})$ represent the predicted macro-states obtained by encoding the predictions of micro-states. The mathematical problems that each framework aims to solve are also illustrated in the figure. (c) The various output forms of NIS+, which include the learned macro-dynamics, captured emergent patterns, functions of coarse-graining, and the degree of causal emergence.
@@
Additionally, to apply NIS+ to the data generated by the complex systems of multi-agent systems and cellular automata, the structures of encoder can be extended by defining two kinds of combinations: one is to stack the basic encoder, the other is to concatenate the basic encoder. Here, a basic encoder is the composition of an invertible neural network and a projection operation. By inverting the directions of all the basic encoder, and supplements the empty inputs with Gaussian noise vectors, we can obtain the corresponding decoder.
Numerical experiments demonstrate the effectiveness of NIS+ in capturing emergent behaviors and finding causal emergence in various simulations, including SIR dynamics, bird flocking (Boid model), and the Game of Life. Additionally, NIS+ is applied to find emergent features in real brain data from 830 subjects while watching the same movie. Experiments also demonstrate that the learned dynamical model exhibits improved performance in terms of out-of-distribution generalization compared to other models.
4.2 Implications in Causal Representation Learning and Out-of Distribution Generalization
We found that, as a framework of machine learning, neural information squeezer and plus can be regarded as a kind of framework for causal feature learning Chalupka et al. (2017); Spirtes et al. (2000); Chalupka et al. (2015) or causal representation learning Schölkopf et al. (2021). Causal feature learning aims that to find a high-level causal relations from the micro-level data. In causal representation learning the macro states are causal variables or causal representations Schölkopf et al. (2021), macro-dynamics are causal mechanisms, and the learned coarse-graining strategy is the representation or encoding process. In this sub-section, we will introduce causal representation learning in details at first, and compare the two emerging fields.
4.2.1 Introduction of Causal Representation Learning
Causal representation learning is an emerging field in artificial intelligence, it tries to combine two important fields in machine learning: representation learning and causal inference, together.
Representation learning aims to extract important features (or representations) hidden in data to make predictions automatically Bengio et al. (2013). Deep neural networks are always adopted, important features in data are represented by the activation modes of neurons LeCun et al. (2015); Goodfellow et al. (2016). The modes in different layers usually represent significant features in various scales, and deeper layers always correspond to larger scales Alzubaidi et al. (2021). The features in different layers are combined together to make predictions or classifications. For example, a deep neural network for human face recognition may have two components. The first one is a deep neural network with stacked convolution units LeCun et al. (1998), and it tries to extract features of faces on different scales. The second component is always a common feed-forward neural network which can accept the input features of faces and output a label to identify the character represented by the input image . The first component is for representation, and the second one is for classification. The neural network as a whole must learn to extract important features from images to make high precision classification.
Although representation learning has achieved remarkable success in various domains such as image classification LeCun et al. (2015), face recognition Wang et al. (2018), language understanding Mikolov et al. (2013); Devlin et al. (2018); Vaswani et al. (2017), and game playing Mnih et al. (2015); Dabney et al. (2020), it suffers from a critical limitation: it can only capture the information or associations within the data, but not the underlying causal relationships. According to the theory of causal hierarchy by Judea Pearl Pearl et al. (2000); Pearl (2009); Cui and Athey (2022), causality is very different from association because the former always represents a more stable relationship which is invariant in different environments, while the latter may be more dependent on contexts and reflect the limitation of the data. For instance, in the example of face recognition mentioned previously, if all the images in the training data have outdoor backgrounds, e.g. grasslands, lakes, trees, etc., the neural network may not work well on the images with indoor backgrounds Cui and Athey (2022); Zhang et al. (2021). This is because the neural network may learn the fake association between backgrounds but not foregrounds and labels. However, the association is only the reflection of the bias in the training data, but not the real causal relationship.
In causal inference, people use graphs to represent causal relationships. Nodes represent causal variables, and edges which are always combined with other information like structural equations or conditional probabilities to represent causal relations Pearl et al. (2000). In conventional causal inference, while the parameters in a causal model can be adjusted based on data, all the variables and causal relationships are typically extracted manually rather than directly from data itself.

### 25 Feb 2024 v3

@@ 4.1.1 Neural Information Squeezer (NIS)
This architectural design with an invertible neural network not only reduces the complexity of computation but also separates the input information from the micro-states into noise and EI for prediction. The authors proved a mathematical theorem stating that, as the neural network converges through training, the mutual information between adjacent macro-states tends to approach the mutual information of two adjacent-step micro-states in the data. This means that all the information discarded by the encoder is almost irrelevant to the prediction. All the useful information converges into the information bottleneck within the macro-level dynamics learner. Another theorem demonstrates that the degree of information compression for micro-states increases as we reduce the dimension of the macro-states. Thus, the macro-state dynamics learner, acting as a narrower information bottleneck, progressively discards more and more extraneous information from the original micro-state data.
To address the problem of causal emergence identification, we first optimize the neural networks, and the dimension-averaged EI, $\mathcal{J}_{q}$, can be obtained for different $q$. When $q=p$, the derived $\mathcal{J}_{p}$ is the dimension-averaged EI for the micro-dynamics. Finally, the calculation of
@@
is the measure of dimension-averaged causal emergence (see Equation (28). If $\Delta\mathcal{J}>0$, we can state that causal emergence occurs.
Numeric examples are provided to show the effectiveness of the whole framework. Interestingly, the Neural Information Squeezer can handle time-series data with discrete micro-states, even though the whole framework was initially designed for continuous dynamical systems. Particularly, the model determines the same coarse-graining strategy and macro-level dynamic as illustrated in the example of a boolean network with four nodes in Hoel’s original paper, although no information about node grouping and state mapping was provided in the framework.
It is also interesting to compare the NIS with the frameworks and models mentioned in previous sections. Compared to the theory of computational mechanics, the NIS can be treated as a kind of $\epsilon$-machine because any encoded macro-states can be understood as states. When the whole framework is well trained, such that it can precisely predict future micro-states, the encoded macro-states converge to effective states, which can be treated as causal states in computational mechanics. However, the objective that maximizes the effective information (EI) in the NIS has no correspondence.
The NIS also shares a similarity with G-emergence because it adopts the ideas of Granger causality: the effective macro-state is optimized by predicting the micro-state in the next time step. However, there are several obvious differences between the two frameworks. In the theory of G-emergence, the macro-state must initially be selected manually, although it is optimized automatically. In addition, the NIS uses neural networks to predict future states, whereas G-emergence uses auto-regressive techniques to fit the data. Due to the universal approximate theorem (see Hornik (1991); Teshima et al. (2020)), neural networks are more advantageous.
The NIS provides a solution for causal emergence identification from data, and it is the first solution for addressing the problem of the automatic construction of the coarse-graining strategy and macro-dynamics from data for causal emergence identification.
Another important benefit of the NIS is that it essentially addresses a generalized causal emergence identification problem, extending beyond mere justification. This is because following training, the NIS not only identifies whether causal emergence occurs but also obtains a neural network for coarse-graining the data and another for simulating the macro-dynamics. This can provide us with richer information compared to a mere justification.
@@ 4.1.2 Neural Information Squeezer Plus (NIS+)
One of the biggest problems of the NIS is that it does not directly address the EI maximization problem by optimizing the neural networks (encoder, decoder, and dynamics learner). Instead, it searches for the optimal dimension of the macro-state space $q$ such that EI can be maximized. Therefore, for a given $q$, the neural networks are optimal for reconstructing the micro-dynamics (minimizing Equation (46)) but not for EI. In addition, the NIS cannot support complex computations such as coarse-graining functions involving multiple steps of information conversion and discarding.
To address these problems, the NIS+ was proposed, as discussed in Yang et al. (2023). Mathematically, the problem of EI maximization can be transformed into a machine learning problem based on the definition of EI and the variational inequality and probability reweighting technique. After this conversion, the minimization problem can be solved by training three neural networks $\psi_{\omega},f_{\theta},$ and $g_{\theta^{\prime}}$ with parameters $\omega,\theta$, and $\theta^{\prime}$, respectively. Formally, the minimization problem without constraints can be written as:
@@
where $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t}))$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t+1}))$ are the macro-states. $\lambda$ is a Lagrangian multiplier, which is regarded as a hyperparameter in experiments. $w(\boldsymbol{x}_{t})$ is the inverse probability weight, which is defined as:
@@
where $\tilde{p}$ is the new distribution of macro-states $\boldsymbol{y}_{t}$ after intervention for $do(\boldsymbol{y}_{t}\sim U_{q})$, and $p$ is the natural distribution of the data. In practice, $p(\boldsymbol{y}_{t})$ is estimated through kernel density estimation (KDE) Rosenblatt (1956). The approximated distribution, $\tilde{p}(\boldsymbol{y}_{t})$, is assumed to be a uniform distribution, which is characterized by a constant value. Consequently, the weight $w$ is computed as the ratio of these two distributions. Theorems have been proven to ensure the correctness of this conversion. The overall framework for the NIS+ is depicted in Figure 15.
@@
Additionally, to apply the NIS+ to data generated by complex systems like multi-agent systems and cellular automata, the structures of the encoder can be extended by defining two kinds of combinations: stacking the basic encoder and concatenating the basic encoder. Here, a basic encoder is the composition of an invertible neural network and a projection operation. By inverting the directions of all the basic encoders and supplementing empty inputs with Gaussian noise vectors, we can obtain the corresponding decoder.
Serious numerical experiments have demonstrated that the NIS+ can automatically learn emergent dynamics and the coarse-graining strategy from data and can integrate these results to identify causal emergence.
For instance, in experiments utilizing training data generated by introducing Gaussian noise to SIR dynamics, the NIS+ demonstrated superior performance compared to the NIS in recovering the vector field of derivatives that aligned with the actual macro-level SIR dynamics. To assess the ability of the NIS+ to learn emergent dynamics, the authors generated data based on macro-states using the SIR dynamics as the ground truth. However, the NIS+ cannot directly utilize this data; only micro-level noisy data can be employed. These micro-level data were generated by introducing Gaussian noise to the macro-states.
In the experiments conducted with the Boid model, training data were generated using the renowned Boid model, which simulates the herding behavior of birds. Both the NIS+ and NIS demonstrated the ability to automatically learn macro-level states and dynamics. In this simulation, the authors divided all artificial birds (boids) into two distinct groups. Interestingly, the learned macro-states could also be categorized into two groups. Within each group, the predicted trajectories generated by the learned macro-dynamics of this group accurately followed the center of the boid groups.
Figure 15: The workflow and architecture of our proposed framework, the Neural Information Squeezer Plus (NIS+), for discovering causal emergence within data. (a) Various forms of input data. (b) The framework for our proposed model, the NIS+, incorporates our previous model, the NIS. The boxes represent functions or neural networks, and an arrow pointing to a cross represents the operation of information discarding. $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t+1}$ represent the observational data of micro-states, whereas $\hat{\boldsymbol{x}}_{t+1}$ represents the predicted micro-state. $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})$ represent the macro-states obtained by encoding the micro-states using the encoder. $\hat{\boldsymbol{y}}_{t}=\phi(\hat{\boldsymbol{x}}_{t})$ and $\hat{\boldsymbol{y}}_{t+1}=\phi(\hat{\boldsymbol{x}}_{t+1})$ represent the predicted macro-states obtained by encoding the predictions of micro-states. The mathematical problems that each framework aims to solve are also illustrated in the figure. (c) The various output forms of the NIS+, which include the learned macro-dynamics, captured emergent patterns, functions of coarse-graining, and the degree of causal emergence.
The ability to capture emergent patterns was verified through experiments conducted in the “Game of Life”. The NIS+ demonstrated the capability to automatically discover both static and dynamic patterns, such as the “glider”, within the learned latent space. Additionally, when clear emergent patterns appeared in the data, larger effective information (EI) was observed for the learned emergent macro-dynamics.
Across all the training data generated through simulations, the NIS+ outperformed comparative models, including the NIS, variational autoencoders, and feed-forward neural networks, in experiments testing the model on data distributions that differed from the training data.
The paper also presented results on real data. Two sets of real fMRI time-series data from human subjects were used. The first dataset, AOMIC ID1000 Snoek et al. (2021), contained fMRI scanning data collected while subjects watched the same movie clip. As a comparison, another fMRI dataset, AOMIC PIOP2 Snoek et al. (2021), containing resting-state data from 50 subjects, was also utilized. The primary training objective was to predict the fMRI signal at the next time step using the learned macro-dynamics and decoder while simultaneously maximizing the effective information (EI) of the macro-level dynamics. Ultimately, a one-dimensional macro-state was found to represent all 100-dimensional micro-state data. Through attribution analysis, the authors found that micro-level signals with larger weights contributing to the one-dimensional macro-state were located in areas associated with visual tasks. However, in the comparison dataset with resting-state data, seven-dimensional macro-states were found to represent the micro-states, and attribution experiments revealed that micro-states with larger weights were distributed across different brain areas.
After the macro-dynamics and coarse-graining strategy were learned, the EIs at both the macro-level and micro-level could be calculated so that causal emergence could be quantified. The experiments showed that the measure of causal emergence depends on different types of noise. If noise was added to the observational data (micro-state), the degree of causal emergence increased with the level of noise. In contrast, if the noise was from the dynamical mechanism, i.e., added to macro-states, the degree of causal emergence decreased with the level of noise. This implies that coarse-graining can learn to eliminate noise in raw data so that better macro-dynamics with larger EI can emerge. However, if the noise is from the internal dynamics, it cannot be removed through coarse-graining.
4.2 Why is Causal Emergence “Causal”?
In the previous sections, we discussed two frameworks that utilize machine learning and neural network techniques to address the challenge of causal emergence identification. Both frameworks place significant emphasis on maximizing EI to learn the coarse-graining strategy and macro-dynamics. Notably, the NIS+ framework extends the NIS approach to achieve real EI maximization, whereas the NIS relies on adjusting the hyperparameter $q$ to optimize EI.
Moreover, the NIS+ has demonstrated an additional capability to generalize data distributions that differ from the training data. This suggests that the NIS+ can learn an invariant causal mechanism that is independent of the input distribution shifts through EI maximization. However, before extending this method to other machine learning scenarios, an essential question needs to be addressed: Is there a genuine relationship between causal emergence, EI maximization, and causality itself?
In the previous sections, we introduced two theoretical frameworks of causal emergence: Hoel’s framework and Rosas’s framework. In both frameworks, the measurement of causal emergence heavily relies on mutual information. However, it is crucial to acknowledge that mutual information primarily quantifies correlation rather than causality. So why do the authors assert that their theories pertain to “causal” emergence?
The key point lies in the research objectives of these frameworks, which focus on a Markov dynamical system. In this context, the dynamical mechanism is assumed to be known, and confounding variables are not considered. Under this assumption, measures related to correlation can effectively indicate causality. While mutual information itself is not a direct measure of causality, it can serve as a proxy for causality when the underlying dynamics are well understood and confounding factors are absent. Therefore, in the context of these frameworks, the use of mutual information as a measure for causal emergence is justified because it captures the relationships between variables in a way that aligns with the assumed known dynamics of the system.
When comparing the two frameworks, Hoel’s framework can be considered more “causal” than Rosas’s framework. This is because Hoel’s framework introduces the do operator in the definition of effective information (EI), as expressed in Equation (7). The do operator allows for interventions and represents a causal manipulation of variables, which strengthens the causal interpretation of the framework. Furthermore, the significance of EI lies in its deep connections with other measures of causation proposed in various fields Comolatti and Hoel (2022). According to the information in Section 3.5, these measures aim to quantify the causal impact of an intervention or treatment on an outcome. By incorporating the do operator, EI aligns with these causal effect measures and provides a valuable tool for assessing causal relationships.
Furthermore, let us delve into the application background of the NIS and NIS+ to understand why these frameworks, based on EI maximization, are considered causal. Firstly, the input data in both frameworks, denoted as $x_{t}$, are generated by a Markovian dynamical system, with no other confounding factors considered, apart from the system itself. As a result, the state of two consecutive time steps, represented as $x_{t}$ and $x_{t+1}$, without loss of generality, can be modeled by a simple causal diagram $x_{t}\rightarrow x_{t+1}$ due to the Markovian property. This causal diagram naturally adheres to the exogeneity assumption in causal inference Pearl (2009), indicating the absence of any confounders between the two states. Consequently, the “do” operator in EI can be replaced with conditional probability, which can be estimated from observations.
Taking all these aspects into consideration, we conclude that the theoretical frameworks of causal emergence and the machine frameworks for causal emergence identification, namely the NIS and NIS+, are indeed related to causality. Moreover, they can be extended to other scenarios within the realm of machine learning. We further explore and discuss this point in the subsequent sections.
4.3 Causal Emergence for Machine Learning
In this sub-section, we introduce how causal emergence can enhance machine learning in out-of-distribution scenarios. It turns out that the do intervention introduced in EI captures the causal dependency from the data generation process and thus complements correlation-based machine learning algorithms.
4.3.1 Out-of-Distribution Generalization
As demonstrated in the previous subsection, the measurement of causal emergence by vanilla effective information Hoel et al. (2013) is defined for stationary dynamics systems in a self-supervised fashion since the input (cause) space and the output (effect) space share the support. However, Hoel generalized effective information to general input-output systems Chvykov and Hoel (2020) (see Section 3.5.5), where the cause and the effect space take arbitrary support, which could either be discrete or continuous. Empowered by the general notation of EI, causal emergence could be applied in supervised machine learning to evaluate the strength of causation between the feature space $\mathcal{X}$ (or its learned representation) and the target space $\mathcal{Y}$, thus enhancing the prediction from the cause (feature) to the effect (target). It is worth noting that a direct fitting from $\mathcal{X}$ to $\mathcal{Y}$ on the observations suffices for common prediction tasks with an i.i.d. assumption, implying that the training and test data are independently and identically distributed.
However, if samples are drawn from outside the training distribution, it is essential to learn a representation space that generalizes from the training to the test environment, known as out-of-distribution (OOD) generalization. Due to the common belief that causal relations generalize better than statistical correlations Arjovsky et al. (2019), causal emergence theory could serve as a criterion for the causation embedded in the representation space. The occurrence of causal emergence indicates the revelation of potential causal factors of the target, thus producing a robust representation space with regard to out-of-distribution generalization.
Introduction to OOD
Before the detailed discussion of the connection between causal emergence and OOD generalization, we begin with a brief introduction to the latter topic. Out-of-distribution generalization toward arbitrary distributional shifts is generally considered impossible Ye et al. (2021). Therefore, it is commonly assumed that there exists an underlying causal mechanism governing the data generation process, which remains invariant to distributional shifts. Formally, according to Pearl’s Structural Equation Model Pearl (2009), which causally formulates data generation, the target $Y$ is supposed to be generated from a part of the feature $X$, i.e., $Y=f(X_{c},\epsilon)$, where $X_{c}$ represents the selected features or a representation of $X$, and $\epsilon$ is a random noise independent of $(X,Y)$. Obviously, $X_{c}$ is the true cause of the target, and the causal mechanism $f$ is presumed to be invariant across training and test environments. On the other hand, the raw feature $X$ might comprise a mixture of causal variables and non-causal variables. The latter could arise from confounding, selection bias, or anti-causal mechanisms, all of which introduce spurious correlations between the target and non-causal variables. The strength and direction of these spurious correlations vary across environments, resulting in distributional shifts that hamper the out-of-distribution performance of machine learning techniques. As a remedy, a prominent branch of OOD algorithms Arjovsky et al. (2019); Huh and Baidya (2022); Zhao et al. (2019); Ganin et al. (2016) learns an invariant representation space distilling the causal variables from the raw feature space. Typically, a metric of the representation is devised to verify the stability of the correlation between the learned representation space and the target and is adopted as a regularization for the optimization objective to encourage the representation learner to extract causal variables from features. The paradigm for invariant representation learning can be expressed as Wang et al. (2022):
where $\phi(X)$ is the learned representation, $f$ is the learned predictor, $\mathcal{R}$ is the task-specific risk function, and $\ell_{reg}$ is the regularization.
Representation learning-based OOD algorithms are diversified in their constraints on the representation space characterized by the regularizers. Each metric of the representation fits a specific data generation pattern and does not guarantee to filter out the causal variables when the setting is violated. However, causal emergence potentially offers a united metric of representation for OOD generalization rooted in causality theory.
Before diving into the application of causal emergence, we first discuss the representative regularizers of representation learning and their limitations. Domain-invariant representation learning proposes an aligned feature space with the constraint $g(X)\perp\mathcal{E}$, where $\mathcal{E}$ denotes the environment index for training and test environments. The regularization is implemented by the metric of Maximum Mean Discrepancy (MMD) distance of representation across environments Li et al. (2018) or an adversarial classifier for the environment index Ganin et al. (2016). However, the alignment constraint provides a generalization guarantee only for a covariate shift Zhao et al. (2019), where the distribution shift is restricted to the features $X$ without perturbing $P(Y|X)$. On the other hand, invariant risk minimization Arjovsky et al. (2019); Koyama and Yamaguchi (2020) pursues an invariant conditional distribution $P(Y|g(X))$ of the representation across environments, which translates to $Y\perp\mathcal{E}\;|\;g(X)$. Invariant risk minimization has been proven to recover the causal variable in a data generation process where the non-causal variable is anti-causally generated by the target but is reported to fail under a covariate shift Ahuja et al. (2021). Further, conditional invariant representation Li et al. (2018); Huh and Baidya (2022) is designed specifically for image classification where the target is modeled as the cause of the feature instead of the effect. The conditional independence $g(X)\perp\mathcal{E}\;|\;Y$ is enforced on the representation space.
It has been theoretically proven that any statistical independence regularization is valid for a subgroup of Structural Equation Models, and it does not generalize across all data generation processes Kaur et al. (2022). As a result, the metric of representation manages to capture the causation between the representation space and the target variable only if a specific causal graph of data generation is given, hindering the general applicability of invariant representation learning. Worse still, the existence of the environment index $\mathcal{E}$ in the independence regularizers introduces the necessity of multiple environments for training procedures, increasing the data collection and labeling expenses. To this end, there is active demand for a unified and general metric of representation for OOD generalization, for which causal emergence might be a good candidate.
Causal Emergence and OOD
The link between OOD generalization and causal emergence can be attributed to the do intervention of effective information. By definition Pearl (2009), intervening in the feature distribution with $do(X)$ modifies the structural equations of all the variables in the feature space to a constant value, whereas the equation for the target remains intact as $Y=f(X_{c})$. On one hand, the assumption of the causal mechanism’s invariance to distributional shift is respected since the causal correlation between causal variables and the target is reserved. On the other hand, the spurious correlation between non-causal variables and the target is removed by the intervention. For example, a non-causal variable anti-causally correlated with the target leads to a structural equation as $X_{n}=h(Y,\epsilon)$, and $do(X_{n})$ replaces the equation with $X_{n}=U$, where $U\sim U_{max}$, and $U_{max}$ represents the maximum entropy distribution on the domain of $X_{n}$. As a result, $X_{n}$ is no longer caused by $Y$, and the spurious correlation vanishes. In another example, if a non-causal variable is correlated with the target by a confounder $E$ (probably in the feature space), the structural equation takes the form of $X_{n}=\phi(E,\epsilon)$ and $Y=f(X_{c},E,\epsilon)$. Then, the do intervention would modify the equation to $X_{n}=U$ and confounding ceases to exist. In summary, the intervention introduced by EI captures the causal correlation between features and the target while suppressing the spurious correlation. It renders EI an ideal metric to assess the causality contained in the representation space and, thus, the generalization of the induced algorithm.
Recall that an increment in effective information implies the occurrence of causal emergence. Therefore, we propose a conjecture that out-of-distribution generalization could be achieved simultaneously with maximizing EI (or normalized EI) for representation learning. Notably, Shannon’s mutual information holds the property that any representation has no gain of mutual information with the target over the raw feature space, i.e., $I(g(X);Y)\leq I(X;Y)$, as per the data processing theorem. Thus, direct fitting inclines to absorb as much information with the target as possible into the representation space. In contrast, EI (or normalized EI) would possibly peak at a medium stage of abstraction of raw features, coinciding with the philosophy of OOD generalization that less could be more. Ideally, at the peak of EI (or normalized EI) where causal emergence occurs, all non-causal features are excluded, and the causal features are revealed, resulting in the most informative representation while remaining invariant to distributional shifts.
The removal of spurious correlations and recovery of causal mechanisms by manipulating feature distributions are principles shared by the reweighting technique widely adopted in the OOD literature. For instance, stable learning Shen et al. (2020a, b); Zhang et al. (2021) is designed for scenarios where collinearity exists in input variables, causing spurious correlations between non-causal variables and the target. It achieves this by reweighting training samples to decorrelate the features, thus reducing collinearity and eliminating spurious correlations. Similar to EI, stable learning is free of environment index labeling. Further, both sample reweighting and feature decorrelation share the philosophy of distribution intervention. In this sense, EI could be viewed as an information-theoretic abstraction of reweighting-based debiasing techniques for OOD generalization.
5 Discussion and Perspectives
We have presented several quantitative theoretical frameworks on causal emergence. However, there are numerous unexplored problems and implications that warrant further discussion. In this section, we address four key topics for future research: causal emergence and causal representation learning, ontological and epidemiological causality and emergence, potential applications in complex systems, and understanding complex systems from the perspective of causal emergence.
5.1 Causal Emergence and Causal Representation Learning
Causal representation learning (CRL) is an emerging field in artificial intelligence (AI) Schölkopf et al. (2021) that combines two important fields in AI: representation learning and causal inference. Representation learning aims to extract important features (or representations) hidden in data to make predictions automatically Bengio et al. (2013). It can be regarded as a typical application of deep learning and has achieved remarkable success in various domains such as image classification LeCun et al. (2015), face recognition Wang et al. (2018), language understanding Mikolov et al. (2013); Devlin et al. (2018); Vaswani et al. (2017), and game playing Mnih et al. (2015); Dabney et al. (2020). However, conventional representation learning suffers from a critical limitation: it can only capture the information of associations within the data but not the underlying causal relationships. Therefore, it is important to consider causality in representation learning.